a child fall from the wall of 100 m height after 1 sec superman saw him from the upper end of wall and fly with a constant speed what should be the speed of superman to caught him at the lower point of wall ?
1
Expert's answer
2020-10-07T10:20:00-0400

solution

given data

height of wall (H)=100 m

time taken by the child to reach the lower point of wall

"H=ut+\frac{1}{2}gt^2"


u=o m/sec

so

"t=\sqrt{\frac{2H}{g}}\\value\quad of\space g=10m/s\\t=4.47s"


after 1 sec superman saw him so he has t2 =4.47 - 1=3.47 s to save the child

speed of superman (v)

"v=\frac{H}{t}=\frac{100}{3.47}\\=28.82 m/sec"


so speed of superman should be 28.82 m/s.